How much longer will it take monthly payments of $1,000 to pay off a $100,000 loan if the monthly compounded rate of interest on the loan is 10.5% instead of 9.75%
Answer by stanbon(75887)

How much longer will it take monthly payments of $1,000 to pay off a $100,000 loan if the monthly compounded rate of interest on the loan is 10.5% instead of 9.75%
-------
1000(1+(0.105/12))^(12t) = 100000
(1.00875)^12t = 100
12t = 2/log(1.00875)
12t = 528.6
time = 44 years
--------------
Cheers,
Stan H.
--------------